# sample.bib

### v2

@@ Abstract
In the face of escalating climate change, achieving significant reductions in greenhouse gas emissions from hard-to-abate industrial sectors is imperative. Carbon Capture and Storage (CCS) represents an essential technological advancement to achieve sustainable decarbonization. This manuscript reports on the bilateral CCS network between Norway and Poland, designed and implemented to leverage their capabilities to expedite technology deployment via mutual cooperation and accelerate CCS initiatives targeted to member states’ challenges across Europe, aiming for a meaningful contribution to climate goals. Norway is renowned for its operational acumen demonstrated through landmark projects like Sleipner and Snøhvit, and its forward-looking initiatives, such as the open-source cross-border Northern Lights project, offer advanced infrastructure and expertise. Conversely, Poland, characterized by its coal-dependent economy and the challenge of decarbonizing extensive industrial emissions, presents a significant geological CO2 storage potential, estimated at over 200 gigatonnes. This study delves into the potential synergies derived from collaborative endeavors in academic education, research and development, industrial implementation, regulatory coherence, and public engagement. By underscoring the reciprocal benefits of such partnership, the study underscores the indispensable role of bilateral cooperation in harnessing CCS’s capabilities to meet ambitious climate objectives, paving the way toward a sustainable and low-carbon future. Additionally, it outlines a scalable model for fostering and supporting broader bi- and multi-lateral collaborations, emphasizing the pivotal role of interconnected networks in shaping effective global climate action strategies.
@@ Introduction
CCS encompasses a suite of complementary technologies that capture CO2 emissions at their source, preventing their release into the atmosphere. Once captured, the CO2 is compressed and transported via pipelines, ships, or other means to suitable geological formations for permanent storage. These formations typically include primarily deep saline aquifers and then depleted oil and gas reservoirs suitable for long-term containment and where CO2 can effectively be sequestered [ringrose2021storage]. CCS offers an actionable pathway for decarbonizing existing industrial facilities to extend their operational lifespan while significantly reducing their emissions and environmental footprint. Furthermore, CCS is indispensable in producing low-carbon hydrogen (blue hydrogen) through steam methane reforming of natural gas, whereby the generated CO2 is captured and stored [massarweh2023blue, bauer2022climate]. This low-carbon hydrogen serves a myriad of applications, including transportation, heating, and industrial processes, underpinning a broader transition to a sustainable energy economy.
@@
In contrast, Poland faces the complex challenge of transitioning from a heavily coal-dependent energy infrastructure. Coal currently generates a significant fraction of Poland’s electricity (approximately 70% in recent years), underpinning various industrial processes and resulting in substantial CO2 emissions [pluta2023scenario, gladysz2020techno]. The Bełchatów Power Station, Europe’s largest lignite-fired power plant with a capacity exceeding 5 GW, epitomizes the scale of industrial emissions in Poland [zak2022energy]. Transitioning Poland’s energy sector to align with both national and EU climate commitments is therefore essential. Nevertheless, Poland possesses robust geological CO2 storage potential, estimated at over 200 Gt, concentrated in deep saline aquifers within the Polish Lowlands and depleted hydrocarbon reservoirs in the Carpathian Foredeep [koteras2020assessment, rutters2021state]. This extensive storage capacity, coupled with urgent reduction needs for industrial emissions, positions Poland as a critical partner in a broader European CCS network.
At a relatively nascent stage compared to Norway in terms of CCS development, Poland requires significant investments in infrastructure, technology adaptation, and regulatory framework establishment. To expedite deployment and effectively implement strategies for success, Poland needs to partner with a competent player with extensive experience in diverse aspects of CCS. Through strategic partnerships, Poland can leverage existing expertise, reduce initial barriers to implementation, and position itself as a proactive player in the broader European CCS network, ultimately contributing to national and regional decarbonization goals.
@@ 1 Poland at the Decarbonization Crossroad
Poland finds itself at a significant crossroads as it navigates the complex path of decarbonization, attempting to reconcile its coal-dependent industrial legacy with the EU’s ambitious climate targets. Historically, Poland has been one of the most carbon-intensive economies in Europe [antosiewicz2022distributional], primarily due to its heavy reliance on coal for electricity generation—a stark contrast with the EU average. As of 2021, coal-fired power plants contributed to approximately 70% of Poland’s electricity mix [qvist2020retrofit]. This dependence poses a substantial challenge within the framework of the European Green Deal, which aims for climate neutrality by 2050 and requires a fundamental transformation in how energy is produced and consumed.
@@
Moving forward, incentivizing investments in renewables, streamlining bureaucratic processes, and deploying CCS at industrial sites are crucial measures. Enhanced public awareness and educational initiatives, alongside governmental transparency in climate strategies, will further solidify Poland’s position in meeting EU climate mandates. The international cooperation could catalyze the transition, turning Poland’s decarbonization challenges into opportunities for a sustainable and prosperous future. Poland’s engagement in CCS initiatives and collaboration with Norway may present a transformative potential within its climate strategy. Participating in projects like the EU CCS Interconnector [gravaud2023ccus] can integrate into the broader European CO2 storage infrastructure, ensuring a more resilient and secure emissions containment strategy.
@@ 2.1 Carbon capture and storage potential
Poland’s greenhouse gas emission profiles reveal key contributors [McKinsey2020, owid-CO2-dataset-sources, crippa2024insights, iea2021greenhouse]: power and heat generation accounts for 38% (156 MtCO2e) of emissions, with the majority stemming from coal and lignite combustion. Industry contributes 22% (91 MtCO2e), dominated by subsectors including fuel manufacturing and chemicals. Transportation contributes 15% (63 MtCO2e) and buildings account for 11% (46 MtCO2e), primarily from coal and gas usage for heating. Agriculture, contributing 11% (44 MtCO2e), is characterized by non-CO2 emissions, with waste and miscellaneous sectors adding an additional 3%. Notably, Poland benefits from -34 MtCO2e of negative emissions [EULULUCF], predominantly via carbon removal in the Land Use, Land-Use Change, and Forestry (LULUCF) sector, primarily through existing forests and increased forest coverage.
The geological landscape of Poland promises significant CO2 storage potential [lubon2020co, urych2022numerical], offering diverse formations: saline aquifers possess an estimated capacity of 12,000 million tonnes; depleted hydrocarbon reservoirs add 1,250 million tonnes, with additional offshore deposits in the Polish economic zone contributing another 860 million tonnes. The total estimated static capacity of 14,110 million tonnes of CO2 suggests potential storage for 50-75 years of Poland’s current ETS emissions levels. Saline aquifers, particularly in regions like Bełchatów, Warsaw, and Kujawy, hold the most significant potential. However, detailed assessments are necessary due to potential conflicts and to confirm the capacity and sealing properties in formations such as the Palaeozoic basins.
@@
2.2 Past projects and initiatives
@@
The recent Poland-EU CCS Interconnector at the Port of Gdansk [gravaud2023ccus], which will establish a hub for CO2 import-export utilizing multi-modal transport connectivity with European networks, may accommodate up to 8.7 million tonnes of annual CO2 transport by 2035. The ECO2CEE project solidifies the infrastructure, leveraging multiple transport modes, including rail, road, waterway, and pipeline options for optimized CO2 distribution [gravaud2023ccus, wojnicki2023perspectives]. The Go4ECOPlanet initiative aims to decarbonize Lafarge’s cement processing in Kujawy via advanced CryoCap capture systems, targeting total decarbonization by transporting and storing captured CO2 offshore [strojny2024preliminary].
@@ 2.4 Outlook and strategic steps for deployment
The draft Polish Hydrogen Strategy [PolHyd] offers a glimpse of CCS potential in clean hydrogen production from natural gas by 2030, pending public consultation outcomes. Additionally, the non-binding social agreement with mining unions outlines a gradual coal phase-out by 2049, incorporating CCU/CCS projects like coal gasification with Integrated Gasification Combined Cycle (IGCC) technology and CO2 capture. However, uncertainties in financing and realization timelines remain significant hurdles.
@@
As the industrial and legislative environment evolves, steps to integrate CCS into the National Energy and Climate Plan are crucial for coherence with EU mandates such as the Net Zero Industry Act. This Act mandates significant CO2 injection capabilities by 2030 and introduces streamlined administrative processes, including designated points of contact for CCS permitting to expedite project approvals. Aligning national policies with EU objectives is vital for leveraging financial and regulatory support. Looking forward, Poland should anticipate an expansive role for CCS by 2040, utilizing developed transport networks and regional hubs. Effective implementation of specialized projects like CCS-ready power stations will be pivotal in contributing to Poland’s climate neutrality aspirations. The Ministries of Climate and Environment, alongside the Ministry of Energy, are instrumental in charting this path, supported by industrial stakeholders and research institutions.
@@ 3 Norway’s CCS Endeavors Encapsulated
Norway’s climate targets include a 40% reduction in greenhouse gas emissions by 2030 compared to 1990 levels and climate neutrality by 2050 [NorClim2130, malka2023energy]. The engagement in CCS was reinforced when the integration of EU CCS Directive (2009/31/EC) into Norwegian law. However, Norway’s experience with CCS predates this directive. The Sleipner project, operational since 1996, serves as a key example, safely storing over 20 million tonnes of CO2 to date [furre2019building, furre2024sleipner]. Similarly, the Snøhvit project, initiated in 2008, captures and stores approximately 0.7 million tonnes of CO2 annually [furre2019building, ringrose2020store].
@@
Companies are increasingly interested in storing CO2 on the NCS, with the government looking to support profitable carbon sequestration businesses. Companies with the necessary expertise and clear business plans can apply to the Ministry of Energy for a permit tailored to their needs. By October 2024, the Ministry had granted eleven permits under CO2 storage regulations, ten in the North Sea and one in the Barents Sea. Additionally, in June 2024, three new storage areas were announced, adding to the existing regions, all situated in the North Sea.
There is growing interest among companies in CO2 storage on the NCS, supported by government initiatives aimed at fostering profitable carbon sequestration ventures. Companies possessing the requisite expertise and well-defined business strategies are encouraged to apply to the Ministry of Energy for permits specifically tailored to their operational needs. By October 2024, the Ministry had issued eleven permits under CO2 storage regulations, ten of which are located in the North Sea, and one in the Barents Sea. Furthermore, in June 2024, the announcement of three new storage areas expanded the existing zones, all positioned in the North Sea.
In addition to its substantial storage capacity, Norway offers a robust infrastructure for CCS operations. Existing offshore infrastructure, originally developed for oil and gas activities, presents opportunities for repurposing into CO2 transport systems. The Long Ship project further enhances this capability by establishing dedicated CO2 transport networks, utilizing custom-designed CO2-transport ships and a receiving terminal anticipated to handle 1.5 million tonnes of CO2 annually by 2024. The project’s future phases aim to expand this capacity beyond 5 million tonnes. In this context, Norway is considering CO2 shipping as a viable transportation strategy, thereby augmenting the flexibility of its infrastructure and facilitating international CO2 reception via specialized carriers.
@@ 4.1 Synergic collaboration components
Given the multifaceted dynamics of excellence and challenges inherent at both ends of this cooperation and based on an analysis of near-term and long-term perspectives outlined in previous sections, the overarching aim of this initiative was set to establish a sustainable bilateral network. This network would connect Norwegian and Polish entities with the long-term vision of fostering collaborations across the entire CCS value chain. This involves bridging the industry, business, and research sectors to expedite CCS deployment initiatives. Key institutions such as AGH University of Krakow, the University of Oslo, and Norwegian Energy Partners were integral to this endeavor. The project team has outlined activities with well-defined objectives: aligning needs with expertise and implementation pathways, facilitating direct contact and matchmaking among stakeholders, and promoting competence-building and infrastructure development, with a strong emphasis on proactive political and public engagement. To actualize these goals, several strategic initiatives have been planned:
@@
The synergic collaboration pathways, shown in Figure 2, visualize the designed framework underpinning the collaborative efforts between Norway and Poland. The nodes within the graph represent critical components such as stakeholder alignment, public engagement, and workshops, capturing the fundamental elements guiding this collaboration. By interlinking nodes via weighted edges, the visualization underscores the dynamic interplay and influence each component exerts within the synergistic framework—indicating pathways like knowledge transfer and economic policy support, which collectively drive cohesive action. Clusters are proposed to group related components into thematic domains, such as Stakeholder & Public Engagement and Cross-border Collaboration, which illustrate the broad-based integration of efforts required to meet ambitious decarbonization goals. The Capstone Symposium node, epitomizes the culmination of strategic discussions, reinforcing the collective-driven approach to CCS initiatives. Through these elements, the visualization encapsulates the seamless convergence of business, education, and innovation strategies deployed within this bilateral alliance, thereby providing an insightful overview of how Norway and Poland may foster operational excellence and meet the challenges of today’s low-carbon initiatives.

### v3

@@ Abstract
In the face of escalating climate change, achieving significant reductions in greenhouse gas emissions from hard-to-abate industrial sectors is imperative. Carbon Capture and Storage (CCS) represents an essential technological advancement to achieve sustainable decarbonization. This manuscript reports on the bilateral CCS network between Norway and Poland, designed and implemented to leverage their capabilities to expedite technology deployment via mutual cooperation and accelerate CCS initiatives targeted to member states’ challenges across Europe, aiming for a meaningful contribution to climate goals. Norway is renowned for its operational acumen demonstrated through landmark projects like Sleipner and Snøhvit, and its forward-looking initiatives, such as the open-source cross-border Northern Lights project, offer advanced infrastructure and expertise. Conversely, Poland, characterized by its coal-dependent economy and the challenge of decarbonizing extensive industrial emissions, presents a significant geological CO2 storage potential, estimated at over 15.5 gigatonnes. This study delves into the potential synergies derived from collaborative endeavors in academic education, research and development, industrial implementation, regulatory coherence, and public engagement. By underscoring the reciprocal benefits of such partnership, the study underscores the indispensable role of bilateral cooperation in harnessing CCS’s capabilities to meet the EU’s ambitious climate objectives, paving the way toward a sustainable and low-carbon future. Additionally, it outlines a scalable model for fostering and supporting broader bilateral and multi-lateral collaborations, emphasizing the pivotal role of interconnected networks in shaping effective global climate action strategies.
@@ Introduction
CCS encompasses a suite of complementary technologies that capture CO2 emissions at their source, preventing their release into the atmosphere. Once captured, the CO2 is purified, compressed, and transported via pipelines, ships, or other means to suitable geological formations for permanent storage. These formations typically include primarily deep saline aquifers and then depleted oil and gas reservoirs suitable for long-term containment and where CO2 can effectively be sequestered [ringrose2021storage]. CCS offers an actionable pathway for decarbonizing existing industrial facilities to extend their operational lifespan while significantly reducing their emissions and environmental footprint. Furthermore, CCS is indispensable in producing low-carbon hydrogen (blue hydrogen) through steam methane reforming of natural gas, whereby the generated CO2 is captured and stored [massarweh2023blue, bauer2022climate]. This low-carbon hydrogen serves a myriad of applications, including transportation, heating, and industrial processes, underpinning a broader transition to a sustainable energy economy.
@@
In contrast, Poland faces the complex challenge of transitioning from a heavily coal-dependent energy infrastructure. Coal currently generates a significant fraction of Poland’s electricity (approximately 60% in 2023 as shown in Fig. 1), underpinning various industrial processes and resulting in substantial CO2 emissions [pluta2023scenario, gladysz2020techno]. The Bełchatów Power Station, Europe’s largest lignite-fired power plant with a capacity exceeding 5 GW, epitomizes the scale of industrial emissions in Poland [zak2022energy]. Transitioning Poland’s energy sector to align with both national and EU climate commitments is therefore essential. Nevertheless, Poland possesses robust geological CO2 storage potential, estimated at over 15 Gt [CLAT2025], concentrated in deep saline aquifers within the Polish Lowlands and depleted hydrocarbon reservoirs in the Carpathian Foredeep [koteras2020assessment, rutters2021state]. This extensive storage capacity, coupled with urgent reduction needs for industrial emissions, positions Poland as a critical partner in a broader European CCS network.
Poland’s CCS development is relatively nascent compared to Norway’s. Therefore, Poland requires significant investments in infrastructure, technology adaptation, and regulatory framework establishment. To expedite deployment and effectively implement strategies for success, Poland needs to partner with a competent player with extensive experience in diverse aspects of CCS. Through strategic partnerships, Poland can leverage existing expertise, reduce initial barriers to implementation, and position itself as a proactive player in the broader European CCS network, ultimately contributing to national and regional decarbonization goals.
@@ 1 Poland at the Decarbonization Crossroad
Poland finds itself at a significant crossroads as it navigates the complex path of decarbonization, attempting to reconcile its coal-dependent industrial legacy with the EU’s ambitious climate targets. Historically, Poland has been one of the most carbon-intensive economies in Europe [antosiewicz2022distributional], primarily due to its heavy reliance on coal for electricity generation (Fig. 1)—a stark contrast with the EU average. As of 2024, coal-fired power plants contributed to approximately 60% of Poland’s electricity mix [qvist2020retrofit] (Fig.1), which dropped around 10% in the past couple of years due to large-scale deployment of renewable energy sources. This dependence still poses a substantial challenge within the framework of the European Green Deal, which aims for climate neutrality by 2050 and requires a fundamental transformation in how energy is produced and consumed.
@@
Moving forward, incentivizing investments in renewables, streamlining bureaucratic processes, and deploying CCS at industrial sites are crucial measures. Enhanced public awareness and educational initiatives, alongside governmental transparency in climate strategies, will further solidify Poland’s position in meeting EU climate mandates. The international cooperation could catalyze the transition, turning Poland’s decarbonization challenges into opportunities for a sustainable and prosperous future. Poland’s engagement in CCS initiatives and collaboration with Norway may present a transformative potential within its climate strategy. Participating in projects like the ECO2CEE (a continuation of the EU CCS Interconnector) [gravaud2023ccus, Orlan2025] can integrate into the broader European CO2 storage infrastructure, ensuring a more resilient and secure emissions containment strategy.
@@ 2.1 Carbon capture and storage potential
Poland’s greenhouse gas emission profiles reveal key contributors [McKinsey2020, owid-CO2-dataset-sources, crippa2024insights, iea2021greenhouse] (Fig.1): power and heat generation accounts for 38% (156 MtCO2e) of emissions, with the majority stemming from coal and lignite combustion. Industry contributes 22% (91 MtCO2e), dominated by subsectors including fuel manufacturing and chemicals. Transportation contributes 15% (63 MtCO2e) and buildings account for 11% (46 MtCO2e), primarily from coal and gas usage for heating. Agriculture, contributing 11% (44 MtCO2e), is characterized by non-CO2 emissions, with waste and miscellaneous sectors adding an additional 3%. Notably, Poland benefits from -34 MtCO2e of negative emissions [EULULUCF], predominantly via carbon removal in the Land Use, Land-Use Change, and Forestry (LULUCF) sector, primarily through existing forests and increased forest coverage.
As shown in Figure 2, the geological resources of Poland promise significant CO2 storage potential [lubon2020co, urych2022numerical], offering diverse formations: saline aquifers possess an estimated capacity of 14,3 Giga tonnes (Gt); depleted hydrocarbon reservoirs add 1,0 Gt, potentially 0.1 Gt storage in coalbed fields. There is a suggested possibility of additional offshore pore space in the Polish economic zone contributing another 0.8 Gt. The total estimated static capacity of CO2 sequestration suggests potential storage for the foreseeable future (approximately 50-75 years) of Poland’s current ETS emissions levels. Saline aquifers, particularly in regions like Bełchatów, Warsaw, and Kujawy, hold the most significant potential. However, detailed assessments are necessary due to potential conflicts and to confirm the capacity and sealing properties in formations such as the Palaeozoic basins.
@@
2.2 Past projects and present initiatives
@@
The recent ECO2CEE (former Poland-EU CCS Interconnector) at the Port of Gdansk [gravaud2023ccus, Orlan2025] (Fig.2), which will establish a hub for CO2 import-export utilizing multi-modal transport connectivity with European networks, may accommodate up to 8.7 million tonnes of annual CO2 transport by 2035. The ECO2CEE project solidifies the infrastructure, leveraging multiple transport modes, including rail, road, waterway, and pipeline options for optimized CO2 distribution [gravaud2023ccus, wojnicki2023perspectives]. The Go4ECOPlanet initiative aims to decarbonize Holcim’s cement processing in Kujawy via advanced CryoCap capture systems, targeting total decarbonization by transporting and storing captured CO2 offshore [strojny2024preliminary].
The latest strategy of the ORLEN Group, unveiled in January 2025 [Orlan20252], highlights the significance of CCS and CCU technologies across the downstream, upstream, and supply value chains. By 2035, the proposed carbon management services aim to encompass the terminal in Gdansk and achieve a transport and storage capacity of 4 million tons of CO2 annually. This target includes 1.1 million tons per annum (Mtpa) for ORLEN Group’s own needs, 1 Mtpa allocated to the Go4ECOPlant initiative for external capacity, and 1.9 Mtpa designated for commercial capacities.
@@ 2.4 Outlook and strategic steps for deployment
The Polish Hydrogen Strategy [PolHyd] offers a glimpse of CCS potential in clean hydrogen production from natural gas by 2030, pending public consultation outcomes. Additionally, the non-binding social agreement with mining unions outlines a gradual coal phase-out by 2049, incorporating CCU/CCS projects like coal gasification with Integrated Gasification Combined Cycle (IGCC) technology and CO2 capture. However, uncertainties in financing and realization timelines remain significant hurdles.
@@
As the industrial and legislative environment evolves, steps to integrate CCS into the National Energy and Climate Plan are crucial for coherence with EU mandates such as the Net Zero Industry Act. This Act mandates significant CO2 injection capabilities by 2030 and introduces streamlined administrative processes, including designated points of contact for CCS permitting to expedite project approvals. Aligning national policies with EU objectives is vital for leveraging financial and regulatory support. Looking forward, Poland should anticipate an expansive role for CCS by 2040, utilizing developed transport networks and regional hubs. Effective implementation of specialized projects like CCS deployment in hard-to-abate industry and potential CCS-ready power stations will be pivotal in contributing to Poland’s climate neutrality aspirations. The Ministries of Climate and Environment, alongside other Ministries, are instrumental in charting this path, supported by industrial stakeholders and research institutions.
@@ 3 Norway’s CCS Endeavors Encapsulated
Norway’s climate targets include a 40% reduction in greenhouse gas emissions by 2030 compared to 1990 levels and climate neutrality by 2050 [NorClim2130, malka2023energy]. The engagement in CCS was reinforced when the integration of EU CCS Directive (2009/31/EC) into Norwegian law. However, Norway’s experience with CCS predates this directive. The Sleipner project, operational since 1996, serves as a key example, safely storing over 20 million tonnes of CO2 to date [furre2019building, furre2024sleipner] ((Fig.3)). Similarly, the Snøhvit project, initiated in 2008, captures and stores approximately 0.7 million tonnes of CO2 annually [furre2019building, ringrose2020store].
@@
There is growing interest among companies in CO2 storage on the NCS, supported by government initiatives aimed at fostering profitable carbon sequestration ventures. Companies possessing the requisite expertise and well-defined business strategies are encouraged to apply to the Ministry of Energy for permits specifically tailored to their operational needs, as shown in Figure 3. By October 2024, the Ministry had issued eleven permits under CO2 storage regulations, ten of which are located in the North Sea, and one in the Barents Sea. Furthermore, in June 2024, the announcement of three new storage areas expanded the existing zones, all positioned in the North Sea.
In addition to its substantial storage capacity in NCS, Norway offers a robust infrastructure for CCS operations. Existing offshore infrastructure, originally developed for oil and gas activities, presents opportunities for repurposing into CO2 transport systems. The Long Ship project further enhances this capability by establishing dedicated CO2 transport networks, utilizing custom-designed CO2-transport ships and a receiving terminal anticipated to handle 1.5 million tonnes of CO2 annually by 2024. The project’s future phases aim to expand this capacity beyond 5 million tonnes. In this context, Norway is considering CO2 shipping as a viable transportation strategy, thereby augmenting the flexibility of its infrastructure and facilitating international CO2 reception via specialized carriers.
Figure 3: Map of the Norwegian Continental Shelf (NCS) showing petroleum fields offshore Norway and eight CO2 storage licenses in the North Sea. It also presents active CCS facilities, such as the Sleipner storage site. Note that the Polaris license and the Snøhvit geological storage site in the Barents Sea are not included within the scope of this map. The data and map elements are adapted from [SD2025, Eq2025].
@@ 4.1 Synergic collaboration components
Given the multifaceted dynamics of excellence and challenges inherent at both ends of this cooperation and based on an analysis of near-term and long-term perspectives outlined in previous sections, the overarching aim of this initiative was set to establish a sustainable bilateral network. This network would connect Norwegian and Polish entities with the long-term vision of fostering collaborations across the entire CCS value chain. This involves bridging the industry, business, and research sectors to expedite CCS deployment initiatives. Key institutions such as AGH University of Krakow, the University of Oslo, and Norwegian Energy Partners were integral to this endeavor. The project team has outlined activities with well-defined objectives: aligning needs with expertise and implementation pathways, facilitating direct contact and matchmaking among stakeholders, and promoting competence-building and infrastructure development, with a strong emphasis on proactive political and public engagement. To achieve these goals, a series of strategic initiatives have been outlined, as depicted in Figure 4:
@@
The synergic collaboration pathways (Fig. 4), visualize the designed framework underpinning the collaborative efforts between Norway and Poland. The nodes within the graph represent critical components such as stakeholder alignment, public engagement, and workshops, capturing the fundamental elements guiding this collaboration. By interlinking nodes via weighted edges, the visualization underscores the dynamic interplay and influence each component exerts within the synergistic framework—indicating pathways like knowledge transfer and economic policy support, which collectively drive cohesive action. Clusters are proposed to group related components into thematic domains, such as Stakeholder & Public Engagement and Cross-border Collaboration, which illustrate the broad-based integration of efforts required to meet ambitious decarbonization goals. The Capstone Symposium node, epitomizes the culmination of strategic discussions, reinforcing the collective-driven approach to CCS initiatives. Through these elements, the visualization encapsulates the seamless convergence of business, education, and innovation strategies deployed within this bilateral alliance, thereby providing an insightful overview of how Norway and Poland may foster operational excellence and meet the challenges of today’s low-carbon initiatives.